A bucket containing water of depth of 10cm is kept in a lift which is vertically upwards with an acceleration which is eaqual to 3/2 times of acceleration due to gravity. Then pressure intensity at the bottom of the bucket in kg/sq.cm is...

Grade:10

1 Answers

Ravleen Kaur
6 years ago
Hello student,
Given:
Acceleration of the system = a = 1.5g
Normal Force on the base of bucket , F= m (g + 1.5 g) =2.5 m kg
mass m = density in kg/cm3X A X h (A - area of the base, h – depth)
m = 10-3X A X 10
F = 0.025A
Pressure on the base in kg/cm3= F/A = 0.025kg/cm3

Regards,
IHE